URGNET QUESTION
hey there ... thank you this has been great help..:) The question is-
The sum of 2 numbers is 12. Find the numbers so that their product is a maximum.
Thankyou soo much..

Answer by stanbon(48558)

The sum of 2 numbers is 12. Find the numbers so that their product is a maximum.
-----
Let one of the numbers be "x"
The other number is (12-x).
-------------------
Product = x(12-x)
p(x) = -x^2+12x
Maximum occurs when x = -b/2a = -12/(2*-1) = 6
P(6) = -6^2+12*6 = -36+72 = 36
------------------------
The numbers that give a maximum product are:
x = 6
12-x = 12-6 = 6
==========================
Cheers,
Stan H.